Ariel is 2 years older than twice Juan's age and Felix is 6 years older than Juan. If the sum of their ages is 80, how old is each person?
Answer by stanbon(75887)

Ariel is 2 years older than twice Juan's age and Felix is 6 years older than Juan. If the sum of their ages is 80, how old is each person?
--------------------
Equations:
a = 2j+2
f = j+6
a + f + j = 80
-----
Substitute for a and f and solve for "j":
2j+2 + j+6 + j = 80
4j = 72
j = 18 (Juan's age)
---
Solve for "a":
a = 2j+2
a = 36+2 = 38 (Ariel's age)
----
Solve for "f":
f = j+6
f = 24 (Felix's age)
----
Cheers,
Stan H.
=========